Grade 12Magical Mathematics[Interesting Approach]

In a single cast with 2 dice, the odds against drawing 7 is?

a)1/6

b)1/12

c)5:1

d)1:5

Answer is given as (c)

Plz explain

Chilukuri Sai Kartik
15 Years agoGrade 12

1 Answer

AKASH GOYAL AskiitiansExpert-IITD
ApprovedApproved Tutor Answer15 Years ago

Dear Chilikuri

First calculate cases of getting 7

(1,6),(2,5),(3,4),(4,3),(5,2),(6,1), out of 6*6=36 cases

So add against=(36-6)/6=5:1

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..